OpenStudy (anonymous):

36/(a/10) a) less than 1 b) greater than

OpenStudy (anonymous):

So the equation is:\[\frac{ 36 }{\frac{ a }{ 10 }}\] Well, it reduces to: \[36*\frac{ 10 }{ a }\] Which reduces to:\[\frac{ 360 }{ a }\] So depending on the value of a, it could be either greater or less than one. Do you have anymore information from the problem?